OpenStudy (anonymous):

how do i Find the slope of the line 5x + 4Y = -3

OpenStudy (katrinakaif):

Slope is -5/4

OpenStudy (anonymous):

write the eqn in the form y=mx+b where m is the slope y=-5/4x-3/4 slope=-5/4

OpenStudy (katrinakaif):

Why? Because we neede to change that equation to y=mx + b format. How? First bring 5x to the other side by substracting it giving: +4y = -5y - 3 then divide by 4 giving y= -5x/4 + 3/4 whatever comes before the xis the slope :)